OpenStudy (anonymous):

write two ratios that are equivalent to 3:1 SHOW WORK PLEASE HOW U GOT IT

OpenStudy (phi):

can you write 3:1 as a fraction?

OpenStudy (anonymous):

no only ratios

OpenStudy (phi):

I mean, write 3:1 as a fraction (it will make the work easier)

OpenStudy (anonymous):

okay but the ratios will be like

OpenStudy (anonymous):

do the work in every way

OpenStudy (anonymous):

but the ratios should have two dots in middle like ex:3:1

OpenStudy (phi):

yes. But to do this problem, write 3:1 as a fraction (as the first step)

OpenStudy (anonymous):

okay

OpenStudy (phi):

\[ \frac{3}{1} \] now to find other fractions that have the same value, multiply the top and bottom by the same number. For example, we can multiply top and bottom by 2: \[ \frac{3 \cdot 2}{1 \cdot 2}= \frac{6}{2} \] 6/2 is another ratio that means the same thing as 3/1 we can write it as 6:2 now find another ratio (multiply top and bottom by a different number than 2)

OpenStudy (anonymous):

how many ratios are there could please write them down 6:2 and

OpenStudy (phi):

there are as many ratios (that are the same as 3:1) as numbers you can think of. (infinite number) do you understand: \[ \frac{3 \cdot n}{1 \cdot n}= \] replace the "n" with a number can you do that ?

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alot more better

OpenStudy (phi):

what do you get if you replace "n" with 3 ?

OpenStudy (anonymous):

u get 3

OpenStudy (phi):

like this: \[ \frac{3 \cdot n}{1 \cdot n} \\ \text{ replace n with 3. You get }\\ \frac{3 \cdot 3}{1 \cdot 3}= \]

OpenStudy (phi):

now simplify the top. what is 3 times 3 ? simplify the bottom. what is 1 times 3 ?

OpenStudy (anonymous):

9 and 3

OpenStudy (phi):

yes, but write it 9/3 or 9:3

OpenStudy (anonymous):

okay very helpful

OpenStudy (phi):

let's find one more. what do you get if you replace n with 4 ?

OpenStudy (anonymous):

not sure about it 2

OpenStudy (phi):

\[ \frac{3 \cdot n}{1 \cdot n} \\ \text{ replace n with 4. You get ?}\]

OpenStudy (anonymous):

12:4

OpenStudy (phi):

yes.

OpenStudy (anonymous):

i am posting something new

OpenStudy (phi):

ok